how many gallons of 15 percent saw solution must be mixed with 30 gallons of 25 percent solution to obtain 20 percent solution
Answer by mananth(16946)

--- percent ---------------- Amount
Solution I 25 ---------------- 30 gallons
solution II 15 ---------------- x gallons
Mixture 20 ---------------- 30 + x gallons

25 * 30 + 15 x = 20 ( 30 + x )
750 + 15 x = 600 + 20 x
15 x -20 x = -750 + 600
-5 x = -150
/ -5
x = 30 gallons solution II

m.ananth@hotmail.ca